\begin{thebibliography}{10}

\bibitem{Baddeley}
R.~W. Baddeley.
\newblock Two-arc transitive graphs and twisted wreath products.
\newblock {\em J. Algebraic Combin.}, 2(3):215--237, 1993.

\bibitem{Biggs-book}
N.~Biggs.
\newblock {\em Algebraic graph theory}, volume No. 67 of {\em Cambridge Tracts in Mathematics}.
\newblock Cambridge University Press, London, 1974.

\bibitem{MAGMA}
W.~Bosma, J.~Cannon, and C.~Playoust.
\newblock The {M}agma algebra system. {I}. {T}he user language.
\newblock {\em J. Symbolic Comput.}, 24(3-4):235--265, 1997.

\bibitem{CLWZ:Irr-covers}
J.~Chen, C.~H. Li, C.~X. Wu, and Y.~Z. Zhu.
\newblock Irreducible covers of $s$-arc transitive graphs.
\newblock {\em in preparation}.

\bibitem{CM1}
M.~D.~E. Conder and J.~Ma.
\newblock Arc-transitive abelian regular covers of cubic graphs.
\newblock {\em J. Algebra}, 387:215--242, 2013.

\bibitem{CM2}
M.~D.~E. Conder and J.~Ma.
\newblock Arc-transitive abelian regular covers of the {H}eawood graph.
\newblock {\em J. Algebra}, 387:243--267, 2013.

\bibitem{DXY}
S.~Du, W.~Xu, and G.~Yan.
\newblock 2-arc-transitive regular covers of {$K_{n,n}$} having the covering transformation group {$\mathbb{Z}^2_p$}.
\newblock {\em Combinatorica}, 38(4):803--826, 2018.

\bibitem{FP2}
X.~G. Fang and C.~E. Praeger.
\newblock Finite two-arc transitive graphs admitting a {R}ee simple group.
\newblock {\em Comm. Algebra}, 27(8):3755--3769, 1999.

\bibitem{FP1}
X.~G. Fang and C.~E. Praeger.
\newblock Finite two-arc transitive graphs admitting a {S}uzuki simple group.
\newblock {\em Comm. Algebra}, 27(8):3727--3754, 1999.

\bibitem{penta_graph}
Y.-Q. Feng, D.-W. Yang, and J.-X. Zhou.
\newblock Arc-transitive cyclic and dihedral covers of pentavalent symmetric graphs of order twice a prime.
\newblock {\em Ars Math. Contemp.}, 15(2):499--522, 2018.

\bibitem{Giudici}
M.~Giudici and C.~S.~H. King.
\newblock On edge-primitive 3-arc-transitive graphs.
\newblock {\em J. Combin. Theory Ser. B}, 151:282--306, 2021.

\bibitem{HNP}
A.~Hassani, L.~R. Nochefranca, and C.~E. Praeger.
\newblock Two-arc transitive graphs admitting a two-dimensional projective linear group.
\newblock {\em J. Group Theory}, 2(4):335--353, 1999.

\bibitem{metacyc_cover}
Z.~Huang and J.~Pan.
\newblock On arc-transitive metacyclic covers of graphs with order twice a prime.
\newblock {\em Electron. J. Combin.}, 25(3):Paper No. 3.9, 17, 2018.

\bibitem{IP}
A.~A. Ivanov and C.~E. Praeger.
\newblock On finite affine {$2$}-arc transitive graphs.
\newblock {\em European J. Combin.}, 14(5):421--444, 1993.
\newblock Algebraic combinatorics (Vladimir, 1991).

\bibitem{Li1}
C.~H. Li.
\newblock The finite vertex-primitive and vertex-biprimitive {$s$}-transitive graphs for {$s\geq 4$}.
\newblock {\em Trans. Amer. Math. Soc.}, 353(9):3511--3529, 2001.

\bibitem{LLW}
C.~H. Li, B.~Ling, and C.~X. Wu.
\newblock A new family of quasiprimitive 2-arc-transitive graphs of product action type.
\newblock {\em J. Combin. Theory Ser. A}, 150:152--161, 2017.

\bibitem{PA-1}
C.~H. Li and A.~Seress.
\newblock Constructions of quasiprimitive two-arc transitive graphs of product action type.
\newblock In {\em Finite geometries, groups, and computation}, pages 115--123. Walter de Gruyter, Berlin, 2006.

\bibitem{EP-s}
C.~H. Li and H.~Zhang.
\newblock The finite primitive groups with soluble stabilizers, and the edge-primitive {$s$}-arc transitive graphs.
\newblock {\em Proc. Lond. Math. Soc. (3)}, 103(3):441--472, 2011.

\bibitem{ZZ}
C.~H. Li, J.~B. Zhang, and H.~Zhou.
\newblock On central covers of arc-transitive graphs.
\newblock {\em preprint}, 2025+.

\bibitem{23-GEN}
M.~W. Liebeck and A.~Shalev.
\newblock Classical groups, probabilistic methods, and the {$(2,3)$}-generation problem.
\newblock {\em Ann. Math. (2)}, 144(1):77--125, 1996.

\bibitem{Lu1}
Z.~P. Lu and R.~Y. Song.
\newblock On basic 2-arc-transitive graphs.
\newblock {\em J. Algebraic Combin.}, 58(4):1081--1093, 2023.

\bibitem{semi_ele_cover}
A.~Malni\v{ c}, D.~Maru\v{ s}i\v{ c}, v.~Miklavi\v{c}, and P.~Poto\v{ c}nik.
\newblock Semisymmetric elementary abelian covers of the {M}\"obius-{K}antor graph.
\newblock {\em Discrete Math.}, 307(17-18):2156--2175, 2007.

\bibitem{Ele_abel_cover}
A.~Malni\v{ c}, D.~Maru\v{ s}i\v{ c}, and P.~Poto\v{ c}nik.
\newblock Elementary abelian covers of graphs.
\newblock {\em J. Algebraic Combin.}, 20(1):71--97, 2004.

\bibitem{ele_cover_pappus}
J.-M. Oh.
\newblock Arc-transitive elementary abelian covers of the {P}appus graph.
\newblock {\em Discrete Math.}, 309(23-24):6590--6611, 2009.

\bibitem{Praeger1}
C.~E. Praeger.
\newblock An {O}'{N}an-{S}cott theorem for finite quasiprimitive permutation groups and an application to {$2$}-arc transitive graphs.
\newblock {\em J. London Math. Soc. (2)}, 47(2):227--239, 1993.

\bibitem{Tutte}
W.~T. Tutte.
\newblock A family of cubical graphs.
\newblock {\em Proc. Cambridge Philos. Soc.}, 43:459--474, 1947.

\bibitem{Weiss}
R.~Weiss.
\newblock The nonexistence of {$8$}-transitive graphs.
\newblock {\em Combinatorica}, 1(3):309--311, 1981.

\bibitem{Wong}
W.~J. Wong.
\newblock Determination of a class of primitive permutation groups.
\newblock {\em Math. Z.}, 99:235--246, 1967.

\bibitem{XDKX}
W.~Xu, S.~Du, J.~H. Kwak, and M.~Xu.
\newblock 2-arc-transitive metacyclic covers of complete graphs.
\newblock {\em J. Combin. Theory Ser. B}, 111:54--74, 2015.

\bibitem{dih_cyclic_cubic_2p}
J.-X. Zhou and Y.-Q. Feng.
\newblock Edge-transitive dihedral or cyclic covers of cubic symmetric graphs of order {$2P$}.
\newblock {\em Combinatorica}, 34(1):115--128, 2014.

\end{thebibliography}

\documentclass[twoside,a4paper,reqno,12pt]{amsart}

\usepackage{cite}
\usepackage{mathrsfs}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage{amssymb,bm}
\usepackage{amsfonts}
\usepackage{longtable}
\usepackage{multirow, hyperref}
\usepackage{braket}
\usepackage{amsmath,enumerate}

\usepackage{graphics}

\def\l{\langle}
\def\r{\rangle}
\def\ov{\overline}

\usepackage{color}
\newcommand{\qd}[1]{{ \fbox{$ \boldsymbol{#1} $} }} 

\def\ZZ{\mathbb Z}
\def\la{\lambda}  \def\m{\mu}
\def\Aut{{\sf Aut}} \def\Cos{{\sf Cos}}
\def\A{{\rm A}} \def\S{\rm S}
\def\PSL{{\rm PSL}}

\def\a{\alpha}
\def\b{\beta}

\def\C{{\bf C}}
\def\N{{\bf N}}
\def\D{{\rm D}}

\newtheorem{thm}{Theorem}[section]

\newtheorem{construction}[thm]{Construction}

\newtheorem{theorem}[thm]{Theorem}
\newtheorem{lemma}[thm]{Lemma}

\newtheorem{corollary}[thm]{Corollary}
\newtheorem{problem}[thm]{Problem}
\newtheorem{example}[thm]{Example}
\theoremstyle{definition}
\newtheorem{rem}[thm]{Remark}

\def\qed{\nopagebreak\hfill{\rule{4pt}{7pt}}
	\medbreak}

\def\Ga{{\it\Gamma}}
\def\Sig{{\it\Sigma}}

\def\Sp{{\rm Sp}}
\def\Sz{{\rm Sz}}

\def\mod{{\sf mod\ }}

\def\S{{\mathrm{S}}}
\def\A{{\mathrm{A}}}

\def\Sym{{\mathrm{Sym}}}
\def\Alt{{\mathrm{Alt}}}

\def\Cos{{\mathrm{Cos}}}

\def\leq{\leqslant}
\def\le{\leqslant}
\def\geq{\geqslant}
\def\ge{\geqslant}

\begin{document}

	\title{Non-solvable $2$-arc-transitive covers of Petersen graphs}

	\author{ Jiyong Chen}
	\author{Cai Heng Li}
	\author{ Ci Xuan Wu}
        \author{ Yan Zhou Zhu}
	
	\address{J.  Chen, School of Mathematics Science \\
		Xiamen University \\
		Xiamen 361005\\
		P. R. China}
	\email{chenjy1988@xmu.edu.cn}
	
	\address{C. H. Li, School of Mathematics Science \\
		Southern University of Science and Technology \\
		Shenzhen 518055\\
		P. R. China}
	\email{lich@sustech.edu.cn}
	
	\address{C. X. Wu, School of Statistics and Mathematics \\
		Yunnan University of Finance and Economics\\
		Kunming 650021\\
		P. R. China}
	\email{wucixuan@gmail.com}

    \address{Y. Z. Zhu, School of Mathematics Science \\
		Southern University of Science and Technology \\
		Shenzhen 518055\\
		P. R. China}
	\email{zhuyz@mail.sustech.edu.cn}
	
	\date\today

\begin{abstract}
We construct connected $2$-arc-transitive covers of the Petersen graph with non-solvable transformation groups, solving the long-standing problem for the existence of such covers.

	\bigskip
	\noindent{\bf Key words:} $2$-arc-transitive, non-solvable cover, minimal cover\\
	\noindent{\bf 2000 Mathematics subject classification:} 05C38, 20B25	
\end{abstract}
		\maketitle

\section{Introduction}

Given a graph $\Gamma=(V,E)$, an \emph{$s$-arc} of $\Gamma$ is a sequence of $s+1$ vertices such that any two consecutive vertices are adjacent and any three consecutive vertices are distinct.
For a group $G$ of automorphisms, a graph $\Gamma$ is said to be \emph{$(G,s)$-arc-transitive} (or simply \emph{$s$-arc-transitive}) if $G$ acts transitively on the set of all $s$-arcs of $\Gamma$.

Finite $s$-arc-transitive graphs form an important class of graphs, and have received considerable attention in the literature.
In particular, from the local action analysis, Weiss \cite{Weiss} proved that there exists no 8-arc-transitive graph except for cycles;
from the global action analysis, Praeger~\cite{Praeger1} initiated a program for the study the class of $2$-arc-transitive graphs, based on taking normal quotient defined below.

Let $\Gamma=(V,E)$ be a $(G,s)$-arc-transitive graph with $s\ge2$.
Let $\mathcal{P}$ be a partition of the vertex set $V$.
Define the \emph{quotient graph} $\Gamma_{\mathcal{P}}$ of $\Gamma$ to be the graph with vertex set $\mathcal{P}$ and two vertices $P_1, P_2$ are adjacent if and only if there exist $u\in P_1$ and $v\in P_2$ which are adjacent in $\Gamma$.
For any normal subgroup $N\lhd G$ which is intransitive on $V$, the $N$-orbits on $V$ form a partition of $V$, and the corresponding quotient graph is called the {\it normal quotient} of $\Gamma$ induced by $N$, denoted by $\Gamma_N$.

It was shown by Praeger~\cite{Praeger1} that if $\Gamma$ is $(G,2)$-arc-transitive and $N$ has at least 3 orbits on $V$, then $\Gamma_N$ is $(G/N,2)$-arc-transitive, and $\Gamma$ and $\Gamma_N$ have the same valency.
In this case, $\Gamma$ is called a {\it cover} of $\Gamma_N$.
To emphasize on the normal subgroup $N$, the graph $\Gamma$ is called a {\it $N$-cover} of $\Gamma_N$, and $N$ is called the {\it transformation group} of this cover.
In other words, the class of 2-arc-transitive graphs is closed under taking normal quotient.
Further, if $\Ga$ is non-bipartite graph and $N\lhd G$ is maximal subject to that $N$ has at least 3 orbits, then each non-identity normal subgroup of $G/N$ is transitive on the vertex set of $\Ga_N$, that is, $G/N$ is {\it quasiprimitive}.

If $\Gamma=(V,E)$ is a $(G,2)$-arc-transitive graph and $G$ is quasiprimitive, then $\Gamma$ has no non-trivial normal quotient, and $\Ga$ is called a quasiprimitive 2-arc-transitive graph.
Based on this observation, Praeger~\cite{Praeger1} established her program to study the class of non-bipartite 2-arc-transitive graphs in two steps:
\begin{enumerate}
	\item characterize quasiprimitive $(G,2)$-arc-transitive graphs;
	\item construct 2-arc-transitive covers of quasiprimitive 2-arc-transitive graphs.
\end{enumerate}

Step~(1) of the program has been fruitful.
Praeger \cite{Praeger1} gave a systematic description about minimal 2-arc-transitive graphs.
In particular, it was shown by Praeger \cite{Praeger1} that if $\Ga=(V,E)$ is $(G,2)$-arc-transitive and $G^V$ is quasiprimitive, then $G$ is of type HA, TW, AS or PA (four out of eight types of quasiprimitive groups).
Further, type~HA is determined in \cite{IP};
type~TW is characterized in \cite{Baddeley};
special cases for type~AS are done in \cite{FP1, FP2,HNP};
examples for type~PA are constructed in \cite{PA-1,LLW}. 
The program has been also successfully applied to characterize some important families of $s$-arc-transitive graphs, see \cite{Giudici,Li1,EP-s,Lu1}.
However, Step~(2) of the program is progressing very slowly, and the existence problem of 2-arc-transitive covers of 2-arc-transitive graphs with non-solvable transformation groups has remained open since then.

There have been efforts in the study of $s$-arc-transitive covers of given $s$-arc-transitive graphs.
J.H. Conway provided a construction of such covers, led to a remarkable result that each connected $s$-arc-transitive graph has a non-trivial connected $s$-arc-transitive cover, refer to \cite[Chapter 19]{Biggs-book}.
However, it is generally hard to construct $s$-arc-transitive $N$-covers $\Gamma$ for a given $(G,s)$-arc-transitive graph $\Sigma$, due to the fact that it is hard to identify a group extension $N.G$ of $N$ by $G$ such that $N.G$ has a connected $s$-arc-transitive graph.
For convenience, if $N$ is a minimal normal subgroup of $N.G$, we say $\Gamma$ to be a {\it minimal normal cover} (or simply a {\it minimal cover}) of $\Sigma$.
Naturally, one should begin with {\it characterizing $s$-arc-transitive minimal covers of minimal $s$-arc-transitive graphs}.

Let $\Sigma=(V,E)$ be a connected minimal $(G,s)$-arc-transitive graph which is not a cycle, and let $\Gamma$ be a connected minimal $s$-arc-transitive cover of $\Sigma$.
Then there exists a group $X=N.G$ such that $N$ is a minimal normal subgroup of $X$ and $G=X/N$ is quasiprimitive on $V$.
Then, by~\cite{ZZ}, one of the following holds: 
\begin{enumerate}[(a)]
\item $N\C_X(N)=N\times M\lhd X$, where $M$ is isomorphic to the socle of $G$;
\item $G$ is faithful on $N$, and $N=T^\ell$ with $T$ simple and $\ell\geq3$.
\end{enumerate}

The problem of characterizing part~(a) is addressed in~\cite{ZZ}, see Example~\ref{ex:diag} for some examples.
In this paper, we focus on case~(b) by studying the following problem.

\begin{problem}\label{prob:cover}
{\rm
Given a connected $(G,s)$-arc-transitive graph $\Sigma$, determine groups $X=N.G$ such that
\begin{enumerate}[(i)] 
\item $N=T^\ell$ with $\ell\geq3$ is the unique minimal normal subgroup of $X$, and
\item there exists a connected $(X,s)$-arc-transitive graph $\Gamma$ which is a $N$-cover of $\Sigma$.
\end{enumerate}
}
\end{problem}

In the literature, there have been lots of publications devoted to the study of covers of $s$-arc-transitive graphs with solvable transformation groups, most of which are with abelian groups of small ranks, refer to 
\cite{DXY, XDKX,CM1,CM2,Ele_abel_cover,semi_ele_cover,dih_cyclic_cubic_2p,penta_graph,metacyc_cover,ele_cover_pappus}. 

In this paper, we construct $2$-arc-transitive covers of the Petersen graph, with non-solvable transformation groups.
For convenience, such covers are called {\it non-solvable covers}.
To the best of our knowledge, these are the first known non-solvable covers of $2$-arc-transitive graphs.




Before stating the main result, we introduce two definitions.
A group $G$ is called \textit{$(2,3)$-generated} if $G=\braket{x,y}$ with $|x|=2$ and $|y|=3$, and called $\textit{perfect}$ if $G=G'$, the commutator subgroup of $G$.

\begin{theorem}\label{Main Theorem}
Let $P$ be a $\{2,3\}$-generated perfect group $P$, and let $G=P\wr\A_5=P^5{:}\A_5$.
There exists a connected $(G,2)$-arc-transitive graph $\Ga$ which is a cover of the Petersen graph.
Moreover, the following properties hold:
	\begin{itemize}
		\item[(i)]$\Aut(\Gamma)=G$;
		\item[(ii)] $\Gamma$ is a non-Cayley graph;
		\item [(iii)] $\Gamma$ has girth $10$.
	\end{itemize}
    
\end{theorem}

\begin{rem}
    In Construction~\ref{cons-1}, we outline the detailed construction of the graph $\Gamma$ presented in  Theorem~\ref{Main Theorem}. This construction need a $(2,3)$-generated perfect group $P$.

\begin{enumerate}[(1)]
\item By Liebeck and Shalev~\cite{23-GEN}, most finite non-abelian simple groups are $(2,3)$-generated, more precisely, except for $\Sz(q)$, $\Sp(4,q)$ with $q$ even, and finitely many other groups, every simple group is $(2,3)$-generated.
See Example~\ref{const-PSL} for concrete examples.

\item If $T$ is $(2,3)$-generated and $N.T=\ZZ_p^\ell.T$ is non-split extension, where $p>3$ and $N$ is minimal normal in $N.T$, then $N.T$ is $(2,3)$-generated.

\item If $T_i$ with $1\leq i\leq d$ are pairwise non-isomorphic simple $\{2,3\}$-groups, then so is the product $T_1\times\dots\times T_d$, as shown in Example~\ref{DirectProductOfSimple}.

\item Construction~\ref{cons-1} can be repeated, namely, if $P$ is perfect and $(2,3)$-generated, then by~Lemma~\ref{cubic-connected}, so are $P\wr\A_5$, $(P\wr\A_5)\wr\A_5$, and so on.
\end{enumerate}
\end{rem}

The following corollary is an immediate consequence of Theorem~\ref{Main Theorem}.

\begin{corollary}
There exist infinitely many connected $2$-arc-transitive non-solvable minimal covers of the Petersen graph.
\end{corollary}

It was shown in \cite{CLWZ:Irr-covers} that, for a prime $p>3$ and a sufficiently large integer $n$ such that $\gcd(p(p-1),n)=1$, the group $P$ discussed in Example~\ref{NonsplitExtension} is a non-split extension of $\ZZ_p^{n^2-1}$ by $\PSL(n,p)$. Moreover, $P$ is a $(2,3)$-generated perfect group.

\begin{corollary}\label{coro-nonsplit}
Let $P=\ZZ_p^{n^2-1}{.}\PSL(n,p)$ be a non-split extension with $p>3$ and $n$ sufficiently large. 
Then there exists a $(P\wr\A_5,2)$-arc-transitive graph which is a cover of the Petersen graph.
\end{corollary}

We remark that for a minimal abelian cover of the Petersen graph the transformation group has rank at most 6, see \cite{CM1}. 
In contrast, the rank of an abelian minimal normal subgroup of the transformation group of the cover given in Corollary~\ref{coro-nonsplit} can be arbitrarily large.

As demonstrated by J.H. Conway's construction~\cite[Chapter 19]{Biggs-book},  each connected $s$-arc-transitive graph has a minimal solvable $s$-arc-transitive cover.
We have shown that  the Petersen graph have minimal non-solvable $2$-arc-transitive covers.
This leads to the following natural problem.
\begin{problem}
    Given a connected $(G,s)$-arc-transitive graph $\Sigma$, does there exist a connected $(X,s)$-arc-transitive graph $\Gamma$ which is a minimal non-solvable cover of $\Sigma$, where $s\geqslant 2$, and $X=T^m.G$ for some non-abelian simple group $T$ and positive integer $m$?
\end{problem}

\section{Basic definitions and some examples}

We collect some definitions and preliminary results in the section, associated with some important examples.

Each arc-transitive graph can be expressed as a coset graph.
Let $\Ga=(V,E)$ be a $G$-arc-transitive graph.
Fix an arc $(\a,\b)$, and an element $g\in G$ which interchanges $\a$ and $\b$, so that $g\in\N_G(G_{\a\b})$.
Since $G$ is transitive on the vertices, $V$ can be identified with the set of right cosets 
$[G:G_\a]=\{G_\a x\mid x\in G\}$ such that the action of $G$ on $V$ is equivalent to the coset action, for any $y\in G$, 
\[y:\ G_\a x\mapsto G_\a xy,\ \mbox{where $x\in G$}.\]
Identify $G_\a$ with $\a$ and $G_\a g$ with $\b$, so that $\Ga(\a)=\{G_\a gh\mid h\in G_\a\}$.
It follows that the adjacency relation is determined by 
\[G_\a x\sim G_\a y\Longleftrightarrow yx^{-1}\in G_\a gG_\a.\]
Then, in this way, $\Ga$ is denoted by $\Cos(G,G_\a, G_\a gG_\a)$, called a coset graph.

A well-known result about $s$-arc-transitive graphs is due to Tutte.

\begin{theorem}{\rm(\cite{Tutte})}\label{Tutte Theorem}
Let $\Ga$ be a cubic $(G,s)$-arc-transitive graph.
Then $s\leq 5$, and $G_\a=\ZZ_3$, $\S_3$, $\S_3\times\S_2$, $\S_4$ or $\S_4\times\S_2$ with $s=1,2,3,4$ or $5$, respectively.
\end{theorem}

We consider the Petersen graph as an example.

\begin{example}\label{ex:Petersen}
{\rm
Let $\Omega=\{1,2,3,4,5\}$, and let $\Omega^{\{2\}}$ be the set of the ten 2-subset of $\Omega$.
Typically, the Petersen graph $\Sig$ is defined as the graph with vertex set $\Omega^{\{2\}}$ such that 
$\a,\b\in\Omega^{\{2\}}$ are adjacent if and only if $\a\cap\b=\emptyset$. 
Fix $\a=\{4,5\}$ and $\b=\{1,2\}$.
The Petersen graph can be expressed as a coset graph.

If we take $G:=\Aut\Sig=\Sym(\Omega)=\S_5$, then $\Sig$ is a $(G,3)$-arc-transitive graph, with stabilizers $G_\a=\Sym\{1,2,3\}\times\Sym\{4,5\}=\S_3\times\S_2$, and $G_{\a\b}=\l(12),(45)\r=\ZZ_2^2$ such that $g=(14)(25)$ interchanges $\a$ and $\b$.
Thus the Petersen graph $\Sig$ can be also written as a coset graph $\Cos(G,G_\a,G_\a gG_\a)$.

On the other hand, if we take $H:=\Alt(\Omega)=\A_5\leq \Aut\Sig$, then $\Sig$ is a $(G,2)$-arc-transitive graph,
with $H_\a=\l (123), (12)(45) \r=\S_3$, and $H_{\a\b}=\l(12)(45)\r=\ZZ_2$ such that $g=(14)(25)$ interchanges $\a$ and $\b$.
Thus the Petersen graph $\Sig$ can be also written as a coset graph $\Cos(H,H_\a,H_\a gH_\a)$.
\qed
}
\end{example}

Conversely, arc-transitive graphs can be constructed as coset graphs from abstract groups.
Let $G$ be a group, $H$ a subgroup of $G$ and $g\in G$. 
Define a \textit{coset graph} $\Ga=\Cos(G,H,HgH)$,
which has vertex set $[G{:}H]$ such that $(Hx, Hy)$ is an arc if and only if $yx^{-1}\in HgH$. 
The following lemma is folklore.
	
\begin{lemma}\label{CosetGraph}
Let $\Ga=\Cos(G, H, HgH)$ be a coset graph, and denote the vertex $H$ by $\a$.
Then $\Ga$ is $G$-arc-transitive and the following statements hold.
\begin{enumerate}[{\rm(i)}]
	\item $\Ga$ is connected if and only if $\braket{H,g}=G$;
	\item $\Ga$ is undirected if and only if $HgH=Hg^{-1}H $;
	\item $\Ga$ is $(G,2)$-arc-transitive if and only if $H$ is $2$-transitive on $[H:H\cap H^g]$.
\end{enumerate}
\end{lemma}

The following example illustrates the method of coset graph construction.

\begin{example}\label{ex:Wong-graph}
{\rm 
Let $p$ be a prime such that $p\equiv \pm1$ $(\mod 16)$, and let $T=\PSL(2,p)$. 
Consider a maximal subgroups $H\cong \S_4$ of $T$  and a subgroup $K\cong \D_8$ of $H$. 
Choose an element $g\in \N_T(K)\setminus H$ such that $g^2 \in K$.
Then $\l H,g\r=T$ as $H$ is a maximal subgroup of $T$, and $H\cap H^{g}=K$.
Thus $\Sig=\Cos(G,H,HgH)$ is a connected cubic $(G,4)$-arc-transitive graph.
This graph was first constructed by Wong in \cite{Wong}. \qed
}
\end{example}

The following example provides a `central' cover with non-solvable transformation groups, which is `trivial' in the sense of group extension.

\begin{example}\label{ex:diag}
{\rm
Take distinct primes $p_1,p_2,\dots,p_n$ such that  $p_i\equiv\pm 1$ $(\mod 16)$.
Let $T_i=\PSL(2,p_i)$.
By Example~\ref{ex:Wong-graph}, there exist $H_i<T_i$ and $g_i\in T_i$ such that $H_i\cong \S_4$, $H_i\cap H_i^{g_i}\cong \D_8$ and $\Ga_i=\Cos(T_i,H_i,H_i g_i H_i)$ is a connected $4$-arc-transitive graph for each $i$. 
In particular, we may assume that $H_i\cap H_i^{g_i}=\langle x_i\rangle{:}\langle y_i\rangle\cong\ZZ_4{:}\ZZ_2$ and $(x_i,y_i)^{g_i}=(x_i,x_iy_i)$.
Let $\sigma_i:H_1\rightarrow H_i$ be isomorphisms such that $(x_1,y_1)^{\sigma_i}=(x_i,y_i)$.
Define $\Ga=\Cos(G,H,HgH)$, where $G=T_1\times T_2\times\dots\times T_n$, $H=\{(h,h^{\sigma_2} ,\dots,h^{\sigma_n})\mid h\in H_1\}\cong\S_4$ and $g=(g_1,g_2,\dots,g_n)$.
Then $H\cap H^g=\langle (x_1,x_2,\dots,x_n)\rangle{:}\langle(y_1,y_2,...,y_n)\rangle \cong \D_8$, and $\Ga=\Cos(G,H,HgH)$ is a 4-arc-transitive graph.
Moreover, the subgroup $\l H,g\r$ projects surjectively to each $T_i$, and so $\l H,g\r=G$ and $\Ga$ is connected.  \qed
}
\end{example}



To end this section, we discuss $(2,3)$-generated perfect groups, which are ingredients for constructing 2-arc-transitive covers of the Petersen graph.
As mentioned before, most finite simple groups are $(2,3)$-generated by Liebeck-Shalev's result.
The following example gives a pair of $(2,3)$-generators for $\PSL(2,p)$.

\begin{example}\label{const-PSL}
{\rm
Let $p \equiv 1 \pmod{4}$ be a prime, and let $T = \PSL(2, p)$. Define elements $a, b \in T$ by
\[
a = \begin{bmatrix} 0 & -1 \\ 1 & -1 \end{bmatrix}, \quad
b = \begin{bmatrix} 0 & 1 \\ -1 & 0 \end{bmatrix},
\]
Computation shows that $a^3 = b^2= 1$. 
Note that the products $ab = \begin{bmatrix} 1 & 0 \\ 1 & 1 \end{bmatrix}$ and $ba = \begin{bmatrix} 1 & -1 \\ 0 & 1 \end{bmatrix}$ both have order $p$. Since no proper subgroup of $\PSL(2, p)$ contains two distinct subgroups of order $p$, it follows that $T=\l a,b\r $. 
\qed }
\end{example}


The next example shows that the direct product of $d$ pairwise non-isomorphic simple $(2,3)$-generated groups is $(2,3)$-generated.

\begin{example}\label{DirectProductOfSimple}
{\rm
Let $d\ge2$ be an integer, and let $T_i$ with $1\le i\le d$ be pairwise non-isomorphic $(2,3)$-generated non-abelian simple groups. More specifically, assume
\[\mbox{$T_i=\l\mu_i,\lambda_i\r$, where $|\mu_i|=2$ and $|\lambda_i|=3$.}\]
Now, let $P=T_1\times \dots \times T_d$ and define $\lambda=(\lambda_1,\dots, \lambda_d)$ and $\mu=(\mu_1,\dots,\mu_d)$. It follows immediately that $|\lambda|=3$ and $|\mu|=2$. Moreover, as the projection of $\l \lambda, \mu\r$ on each of the direct factors is surjective and $T_i$ are pairwise non-isomorphic, we have $P=\l \lambda, \mu\r$ which is a $(2,3)$-generated perfect group.  
\qed
}
\end{example}

The final example of this section provides a method for constructing $\{2,3\}$-generated perfect groups by non-split group extensions.

\begin{example}\label{NonsplitExtension}
{\rm
Assume that $p>3$ is a prime and $\gcd(p(p-1),n)=1$. 
By Liebeck-Shalev's theorem, except for finitely many possible exceptions, $T=\PSL(n,p)$ is $(2,3)$-generated, say $T=\l a,b\r$ where $|a|=3$ and $|b|=2$.
By \cite{CLWZ:Irr-covers}, there exists a non-split extension $P=N.T=\ZZ_p^{n^2-1}.\PSL(n,p)$, where $N=\ZZ_p^{n^2-1}$ is a minimal normal subgroup of $P$.
Let $x,y$ be preimages of $a,b$, respectively, under $P\rightarrow T$.
Then $|x|=3$ or $3p$, and $|y|=2$ or $2p$, and so $|x^p|=3$ and $|y^p|=2$.
Since $N.T$ is a non-split extension and $N$ is a minimal normal subgroup of $P$, it follows that $P$ is perfect and $P=\l x^p,y^p\r$.
\qed
}
\end{example}

\def\calA{{\mathcal A}}

\section{Proof of Theorem~\ref{Main Theorem}}\label{sec:petersencover}

We first present a construction of 2-arc-transitive non-abelian covers of the Petersen graph.

\begin{construction}\label{cons-1}
{\rm
Let $P=\braket{\mu,\la}$ be a perfect group where $\mu^2=\la^3=1$.
Let $G=P\wr \A_5=P^5{:}\A_5$ be such that, for $(t_1,t_2,t_3,t_4,t_5)\in P^5$ and $\pi\in\A_5$,
\[(t_1,t_2,t_3,t_4,t_5)^{\pi^{-1}}=(t_{1^\pi}, t_{2^\pi}, t_{3^\pi}, t_{4^\pi}, t_{5^\pi}).\]
Pick $h_1,h_2,g\in G$ as follows:
\[\begin{array}{l}
	h_1=(1,1,1,\la,\la^{-1})(123),\\
	h_2=(\m,\m,\m,1,1)(12)(45),~\mbox{and}~ \\
	\ \, g=(1,\m,1,1,\m)(14)(25).
\end{array}\]
Let $H=\braket{h_1,h_2}$ and  $\Gamma=\Cos(G,H,HgH)$.
\qed
}
\end{construction}

We shall complete the proof of Theorem~\ref{Main Theorem} by a series of lemmas.

\begin{lemma}\label{lem:covers}
Using the notation defined in Construction~\ref{cons-1}, the following statements hold:
\begin{enumerate}[{\rm(1)}]
\item $H=\langle h_1\rangle{:}\l h_2\r=\S_3$, and $|g|=2$;
\item $H\cap H^g=\l h_2\r=\S_2$;
\item $\Gamma=\Cos(X,H,HgH)$ is a $2$-arc-transitive cubic graph;
\item $\Gamma$ is a normal cover of the Petersen graph.
\end{enumerate}
\end{lemma}
\proof
(1). It is easily shown that $h_1$ is of order 3, and $h_2,g$ are involutions.
Further, $h_1^{h_2}=h_1^{-1}$, and so $H=\l h_1,h_2\r=\S_3$.

(2). Since $h_2g=gh_2$, we obtain that $H\cap H^g=\l h_2\r=\S_2$.

(3). Obviously, $H=\S_3$ acts 2-transitively on $[H:H\cap H^g]$.
Thus $G$ acts 2-arc-transitive on the graph $\Gamma$.

(4). Note that $\overline X=X/N\cong \operatorname{Alt}\{1,2,3,4,5\}=\A_5$.
Let $\overline H$ and $\overline g$ be the homomorphic images of $H$ and $g$, respectively.
Then $\overline H=\l(123),(12)(45)\r=\S_3$, and $\ov g=(14)(25)$.
Then $\Gamma_N=\Cos(\ov X,\ov H,\ov H\ov g\ov H)$ is isomorphic to the Petersen graph by Example~\ref{ex:Petersen}.
\qed

Next, we will prove Theorem~\ref{Main Theorem} by a series of lemmas.

\begin{lemma}\label{ generated by commutators}
The perfect group $P=\l \m,\lambda\r$ is generated by the commutators $[\m,\la]$ and $[\m,\la^2]$.
\end{lemma}
\proof
Let $R=\braket{[\m,\la],[\m,\la^2]}$.
Since $|\mu|=2$,  conjugation by $\mu$ inverts both $[\m,\la]$ and $[\m,\la^2]$, implying that $R^\mu \leq R$.
Since $|\la|=3$,  it follows that $[\m,\la^2]^{\la}=[\la,\mu]=[\mu,\la]^{-1}$.
Note that $[\mu,\la]^\la=[\mu,\la]^{-1}[\mu,\la^2]$.  Consequently, $R^\la \leq R$, and thus $R\unlhd P$.
Since $P/R=\braket{\mu R,\la R}$ is abelian and $P$ is perfect, we conclude that $R=P$.
\qed

Let $G=\braket{h_1,h_2,g}$ . To prove that $\Gamma$ is connected, by Lemma~\ref{CosetGraph}, we only need to show $X=G$. We first state some properties of $G$ in the next lemma.
Write $N=P_1\times\cdots\times P_5$, where $P_i\cong P$ for $1\leq i\leq 5$.
Let $K=G\cap N$.

\begin{lemma}\label{property G}
Let  $S$ be a non-empty subset of $\{1,\dots,5\}$. Then the following statements hold.
\begin{itemize}
\item [(1)] the quotient group $G/K\cong \A_5$, and $G$ is $3$-transitive on $\{P_1,\dots,P_5\}$ by conjugation.
\item[(2)] If $|S|\leq 2$, then $K$ projects surjectively on $\prod_{i\in S}P_i$ for each $S$;
\end{itemize}
\end{lemma}
\begin{proof}
(1). Since $G/K\cong GN/N=\braket{h_1N,h_2N,gN}\cong \braket{(123),(12)(45),(14)(25)}$, it follows that $G/K\cong \A_5$.
Because $\A_5$ is $3$-transitive on $5$ points under the natural action, we have part (1).

(2). We first present some computation results.
Let $x=(h_1g)^5$, $x_1=(\m,1,1,\la,\la^{-1}\m)$ and $x_2=(15234)$  .   Then
 $x=(x_1x_2)^5$ as $h_1g=x_1x_2$. Therefore,
	\begin{equation*}
		\begin{split}
			x&=(x_1x_2)^5=x_1\cdot (x_1)^{x_2^4}\cdot (x_1)^{x_2^3}\cdot (x_1)^{x_2^2}\cdot (x_1)^{x_2}\\
			&=(\m,1,1,\la,\la^{-1}\m)(\la^{-1}\m,1,\la,\m,1)(1,\la,\m,\la^{-1}\m,1)
			(1,\m,\la^{-1}\m,1,\la)(\la,\la^{-1}\m,1,1,\m)\\
			&=(\m\la^{-1}\m\la,\la\m\la^{-1}\m,\la\m\la^{-1}\m,\la\m\la^{-1}\m,\la^{-1}\m\la\m)\\
			&=([\m,\la],[\la^2,\m],[\la^2,\m],[\la^2,\m],[\la,\m]).\\
		\end{split}
	\end{equation*}
	Let $y=(xh_2)^2$. Since $\m$ is an involution, we have $\m[\m,\la^i]\m=[\la^i,\m]$ for $i\ge 0$. Therefore,
	\begin{align}
	y&=(([\m,\la]\m,[\la^2,\m]\m,[\la^2,\m]\m,[\la^2,\m],[\la,\m])(12)(45))^2 \notag\\
			&=([\m,\la]\m,[\la^2,\m]\m,[\la^2,\m]\m,[\la^2,\m],[\la,\m])([\la^2,\m]\m,[\m,\la]\m,[\la^2,\m]\m,[\la,\m],[\la^2,\m])\notag\\
			&=([\m,\la][\la^2,\m]^{\m},[\la^2,\m][\m,\la]^{\m},[\la^2,\m][\la^2,\m]^\m,[\la^2,\m][\la,\m],[\la,\m][\la^2,\m]) \notag\\
			&=([\m,\la][\m,\la^2],[\la^2,\m][\la,\m],1,[\la^2,\m][\la,\m],[\la,\m][\la^2,\m]) \label{eqn-y}
	\end{align}
	Let $z=x^{-1}y$. Then
	$z=([\m,\la^2],[\la,\m],[\m,\la^2],[\la,\m],[\la^2,\m])$.
	
To prove part (2), by the action of $G$ on $\{P_1,\dots,P_5\}$, we only need to show that the projections of $K$ on $P_1$ and $P_1\times P_3$ are surjective.
By Lemma~\ref{ generated by commutators}, the subgroup $\braket{x,y}$ projects surjectively $P_1$, and so does $K$.
Therefore, $K$ projects surjectively on each $P_i$ for $1\le i\le 5$.

Let $\widetilde{K}$ be the projection of $K$ on $P_1 \times P_3$, and let $\pi_1$ and $\pi_3$ denote the projection maps of $\widetilde{K}$ into the first and third direct product, respectively. Then we have $ \pi_1(\widetilde{K}) \cong \pi_3(\widetilde{K}) \cong P $.
Let $\widetilde{P_1}$ and $\widetilde{P_3}$ be the projection of $P_1$ and $P_3$, respectively. 
Note that $\widetilde{K}\cap \widetilde{P}_1 ={\ker \pi_3}=\pi_1(\ker \pi_3)\lhd \pi_1(\widetilde{K})=\widetilde{P}_1$.
As shown in Equation~\ref{eqn-y}, the projection of $y$ is given by $ ([\mu, \lambda][\mu, \lambda^2], 1) $, which lies in $ \widetilde{K} \cap \widetilde{P_1} $. By Lemma~\ref{ generated by commutators}, we have $ P = \langle [\mu, \lambda], [\mu, \lambda][\mu, \lambda^2] \rangle $, implying that $ \widetilde{P_1}/(\widetilde{K} \cap \widetilde{P_1}) $ is cyclic.
Since $P$ is perfect, it follows that $ \widetilde{K} \cap \widetilde{P_1} = \widetilde{P_1}$. Consequently, we conclude that $ \widetilde{K} = \widetilde{P_1} \times \widetilde{P_3}$.
\end{proof}

We fix the following notation to streamline the proofs of the subsequent lemmas.
For any normal subgroup $Q$ of $P$ and $1\leq i\leq 5$, let $Q_i\lhd P_i$ be the corresponding subgroup with $Q_i\cong Q$ such that the conjugation action of $X$ on the set $\{Q_1,\dots,Q_5\}$ is transitive. Denote the direct product $Q_1\times\cdots\times Q_5$ by $Q^5$. 

\begin{lemma}\label{cubic-connected}
The graph $\Gamma$ is connected, and $X=\braket{g,h_1}$ is also a $(2,3)$-generated perfect group. 
\end{lemma}
\begin{proof}
Let $P$ be a minimal counterexample for the statement $X\neq G$. Then for any nontrivial normal subgroup $M$ of $P$, the quotient group $P/M$ is perfect and generated by $\mu M$ and $\la M$.
Let $L=M^5$. Then $X/L\cong (P/M)\wr \A_5$, and by assumption, $X/L=GL/L$. Therefore, $X=GL$, and it follows that $$N=GL\cap N=(G\cap N)L=KM^5.$$

We claim that $Z=Z(P)$, the center of $P$, is trivial.
If $Z\neq 1$, From the discussion above, we have $N=Z^5K$. Thus, $N=N'=(Z^5K)'=K$, and hence $N=K$. Therefore, $X=G$, which contradicts that  $P$ is a counterexample. Hence, our claim holds.

{\bf \noindent Step 1: the projection of $K$ on $P_1\times P_2\times P_3$ is surjective}

Assume the projection of $K$ on $P_1\times P_2\times P_3$ is not surjective.
Let $U=P_1\times P_2\times P_3$ and $V=P_4\times P_5$ and  let $\widetilde{K}$ and $\widetilde{P_i}$ be the projections of $K$ and $P_i$ on $U$, where $1\leq i\leq 3$. Then $\widetilde{K}=KV/V$, $\widetilde{P_i}=P_iV/V$ and $KV<N$.
Let $\pi_{12}$ and $\pi_3$ denote the projection maps of $\widetilde{K}$ into $\widetilde{P}_1\times \widetilde{P}_2$ and  $\widetilde{P}_3$, respectively.
According to Lemma~\ref{property G}(2), $\pi_{12}(\widetilde{K}) =\widetilde{P}_1\times \widetilde{P}_2$ and $\pi_3(\widetilde{K}) =\widetilde{P}_3$.
Since $|\widetilde{P}_1\times \widetilde{P}_2|> |\widetilde{P}_3|$, 
\[|\widetilde{K} \cap (\widetilde{P}_1\times \widetilde{P}_2)| =|\ker \pi_3|>|\ker \pi_{12}| \geq 1.  \]
 This implies that $\widetilde{K}\cap (\widetilde{P_1}\times \widetilde{P_2})\neq 1$.

Suppose $\widetilde{K} \cap \widetilde{P_1} \neq 1$. Since $\widetilde{K}$ projects surjectively onto $\widetilde{P_2} \times \widetilde{P_3}$, it follows that $\widetilde{K} \cap \widetilde{P_1}$ is a nontrivial normal subgroup of $\widetilde{P_1}$.
Thus, there exists a nontrivial normal subgroup $D_1$ of $P_1$ such that $D_1 \leq KV$.
Recall that $D_i\in P_i$ and $D^5=D_1\times\cdots\times D_5$.
Considering that the stabilizer of $G$ on $\{P_4, P_5\}$ is transitive on $\{P_1, P_2, P_3\}$, we deduce that $D_1 \times D_2 \times D_3 \leq KV$.  Then $D^5\leq KV$ as $D_4 \times D_5 \leq V$, and hence $D^5K\leq KV$.
Noting that the assumption has led us to $N =D^5K$, which contradicts $KV<N$.

Therefore, we must have $\widetilde{K} \cap \widetilde{P_1} = 1$, and similarly, $\widetilde{K} \cap \widetilde{P_2} = 1$. Since $$(\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2})) \cap \widetilde{P_1} = \widetilde{K} \cap \widetilde{P_1} = 1,$$
and $\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2}) \lhd \widetilde{P_1} \times \widetilde{P_2}$,
it follows that $\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2})$ centralizes $\widetilde{P_1}$.  Recalling that $Z(\widetilde{P_1})=1$, the centralizer of $\widetilde{P_1}$ in $\widetilde{P_1} \times \widetilde{P_2}$ is $ \widetilde{P_2}$.  Thus $\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2}) \leq \widetilde{P_2}$.  By symmetry, $\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2}) \leq \widetilde{P_1}$.
Consequently, $\widetilde{K} \cap (\widetilde{P_1} \times \widetilde{P_2}) \leq \widetilde{P_1} \cap \widetilde{P_2} = 1$, leading to a contradiction.
Therefore, the projection of $K$ on $P_1 \times P_2 \times P_3$ is surjective.

{\bf \noindent Step 2: $X=G$}

Now, $N=U\times V$ and $K$ projects surjectively on both $U$ and $V$. Then $U/(K\cap U)\cong V/(K\cap V)$, and so $K\cap U\neq 1$ as $|U|>|V|$.
If $K\cap P_1\neq 1$, then $L:=\prod_{i=1}^{5}(K\cap P_i)$ is normal in $X$.
By our assumption, $X/L=G/L$, and hence $X=G$.
Now assume that $K\cap P_i=1$ for each $1\leq i\leq 5$.
Thus $K\cap U$ centralize $P_1$, and so $K\cap U\leq P_2\times P_3$.
Similar arguments for $K\cap P_2$ and $K\cap P_3$, we obtain that
$K\cap U$ is a subgroup of both $ P_1\times P_3$  and $P_1\times P_2$,
Consequently, $K\cap U=1$, which is a contradiction.
Thus $N=K$ and $X=G$, as we required.

By Lemma~\ref{lem:covers}, $h_2$ normalizes $\langle g,h_1\rangle$.
Then $X/\langle g,h_1\rangle\leq \ZZ_2$. As $X=P\wr \A_5$ is perfect, we thus conclude that $X=\langle g,h_1\rangle$.
Therefore, $X$ is a $(2,3)$-genereted group.
\end{proof}

\begin{lemma}\label{N:largestenormal}
Let $M$ be a normal subgroup of  $X$. then either $ M \leq N $ or $M=X$.
In particular, $N$ is a characteristic subgroup of $X$.
\end{lemma}
\begin{proof}
We only need to show that $M=X$ whenever $M\nleq N$. Suppose, for contradiction, that $M\neq N$ and $M<X$. Without loss of generality, assume that $M$ is a maximal normal subgroup of $X$.
Then $ 1 \neq MN/N \trianglelefteq X/N \cong A_5 $. Therefore, $ MN/N = A_5 $
and thus $ X = MN $.  Furthermore, there exist $m\in M$ and $n\in N$ such that $mn=(1 2 3 4 5)$.
If $P_1\not\leq M$, then let $Q_1=P_1\cap M$. For any normal subgroup $R_1$ of $P_1$ satisfying $Q_1< R_1\leq P_1$, we have $M/M \neq R_1M/M  \lhd NM/M=X/M $. By the maximality of $M$, it follows that $R_1M=X$. Therefore,
$$P_1=P_1\cap R_1M=R_1(P_1\cap M)=R_1Q_1=R_1,$$ and hence $Q_1$ is a maximal normal subgroup of $P_1$. Note that $$Q_{i+1}=Q_1^{m^i}\leq M^{m^i}\leq M.$$ It follows that $Q^5$ is a subgroup of $M$. Therefore, $M/Q^5$ is a proper normal subgroup of $X/L$ that is not contained in $N/L$. But $N/L$ is the unique minimal normal subgroup of $X/L\cong (P/Q)\wr \A_5$ as $P/Q$ is non-abelian simple, leading to a contradiction.
If $P_1\leq M$, then $P_{i+1}=P_1^{m^i}\leq M^{m^i}=M$. This gives $N=P^5\leq M$ and $X=MN=M$, a contradiction.

Let $\sigma\in \Aut(X)$. Then $N^\sigma\unlhd X^\sigma=X$. By the above argument, we have $N^\sigma=N$.
Thus $N$ is a characteristic subgroup of $X$.
\end{proof}

For a group $G$ and a subgroup $H$ of $G$, the \textit{core} of $H$ in $G$, denoted ${\sf Core}_G(H)$, is the intersection of all conjugates of $H$ in $G$. It is easy to see $G/{\sf Core}_G(H)$ is a transitive permutation group on the right coset $[G{:}H]$.
The following three lemmas determine the automorphism group of the graph $\Gamma$, its Cayleyness, and its girth.
	\begin{lemma}\label{Aut(Gamma)}
	$\Aut(\Gamma)=X.$
	\end{lemma}
\proof
Let $A=\Aut(\Gamma)$. Denote the vertices $H$ and $Hg$ in $\Gamma$ by $u$ and $ v$, respectively. Then the vertex-, edge-, arc- stabilizers are as follows: $X_u=\l h_1,h_2\r,$ $X_{\{u,v\}}=\l h_2, g\r$ and $X_{(u,v)}=\l h_2\r$.

We first claim that $N\lhd A$.
If  $X\lhd A$, then by Lemma~\ref{N:largestenormal}, $N$ is normal in $A$.
If $\mathsf{core}_A(X)<X$, by Lemma~\ref{N:largestenormal}, we have $\mathsf{core}_A(X) \leq N$, and
 $$
X/\mathsf{core}_A(X) = N/\mathsf{core}_A(X).X/N.
$$
By Theorem~\ref{Tutte Theorem}, $|A{:}X|=|A_u{:}X_u|\leq 8$. Then we have $A/\mathsf{core}_A(X)\leq \S_8$. Since $N$ is a perfect group, $N/\mathsf{core}_A(X)$ is also a perfect group. However, every subgroup of $\S_8$ has at most one non-solvable composition factor. Hence $N/\mathsf{core}_A(X)=1$, and thus $N=\mathsf{core}_A(X)\lhd A$.

By Lemma~\ref{lem:covers}, the quotient graph $\Gamma_N$ is the Petersen graph, so $A/N \leq \S_5$. Suppose $|A{:}X|=2$. Then $A/N \cong \S_5$, and $\Gamma$ is $3$-arc-transitive. By Theorem~\ref{Tutte Theorem}, we deduce that
 $A_{(u,v)} \cong \ZZ_2^2$ and $A_{\{u,v\}} \cong \D_8$ .
Choose $a \in A_{(u,v)} \setminus X_{(u,v)}$. Since $A_{(u,v)}$ is abelian, $h_2^a = h_2$. As $A_{\{u,v\}}$ is non-abelian, $g^a \neq g$. Since $a$ normalizes $X_{\{u,v\}}$, we conclude that $g^a = gh_2$.
Therefore,
$$
\C_N(g)^a = \C_{N^a}(g^a) = \C_N(gh_2).
$$
Let $n=(p_1,p_2,p_3,p_4,p_5)\in N$, where $p_i\in P$ for $1\leq i\leq 5$.
Direct computations give:
\begin{equation*}
	\begin{split}
		ng=&(p_1,p_2\mu,p_3,p_4,p_5 \mu)(14)(25)\\
		gn=&(p_4,\mu p_5, p_3, p_1,\mu p_2)(14)(25).\\
	\end{split}
\end{equation*}
If $n\in \C_N(g)$, then $p_4=p_1$ and $p_5=p_2^\mu$. This gives
\begin{equation}\label{eqn:cng}
    \C_N(g)= \{(p_1, p_2,p_3,p_1,p_2^\mu)\in N | p_1,p_2,p_3\in P\}
\end{equation}
Note that $h_2g=gh_2=(1,\mu,\mu,\mu,1)(15)(24)$.
Similar computations for $n(gh_2)$ and $(gh_2)n$, we obtain that
\begin{equation*}
	\begin{split}
		n(gh_2)=&(p_1,p_2\mu,p_3\mu,p_4\mu,p_5)(15)(24)\\
		(gh_2)n=&(p_5,\mu p_4, \mu p_3, \mu p_2,p_1)(15)(24).\\
	\end{split}
\end{equation*}
If $n\in \C_N(gh_2)$, then $p_5=p_1$, $p_4=p_2^\mu$ and $p_3$ centralizes $\mu$. This gives
\begin{equation}\label{eqn:cngh2}
	\C_N(gh_2)= \{(p_1, p_2,p_3,p_2^\mu,p_1)\in N | p_1,p_2\in P\mbox{ and } p_3\in \C_P(\mu)\}
\end{equation}

Equations~\ref{eqn:cng} and \ref{eqn:cngh2} imply that $\C_N(g)\cong P^3$ and $\C_N(gh_2)\cong P^2\times \C_P(\mu)$.
Since $P=\l \lambda, \mu\r$ is non-abelian, $\C_P(\mu)$ is a proper subgroup of $P$. Consequently, $\C_N(g)\not\cong \C_N(gh_2)$, a contradiction. Hence, we have $A=X$.
\qed

\begin{rem}
    By Lemma~\ref{Aut(Gamma)}, $\Gamma$ is $2$-arc-transitive but not $3$-arc-transitive cover of the Petersen graph. Using Magma, it is not hard to find a subgroup $H\cong \S_3\times \S_2$ and an element $g$ in $\A_5\wr \S_5$ such that the coset graph $\Sigma=\Cos(\A_5\wr \S_5, H, HgH)$ is a $3$-arc-transitive minimal cover of the Petersen graph with transformation group $\A_5^5$. However, it does not seem easy to directly generalize this example to the case of an infinite family.
\end{rem}

\begin{lemma}\label{Non-Cayley}
$\Gamma$ is not a Cayley graph.
\end{lemma}
\begin{proof}
Assume $\Gamma$ is a Cayley graph. By Lemma~\ref{Aut(Gamma)}, $\mathsf{Aut}(\Gamma) = X$ contains a vertex-regular subgroup $K$. Since $\Gamma_N$ is the Petersen graph, $|K| = 10|N|$ and $|X{:}K| = 6$,  so $X/\mathsf{core}_X(K) \leq \S_6$. By Lemma~\ref{N:largestenormal}, we have $\mathsf{core}_X(K) \leq N$.
Now, since $X/\mathsf{core}_X(K) \cong (N/\mathsf{core}_X(K)) . (X/N)$ and $N/\mathsf{core}_X(K)$ is perfect, this forces $N = \mathsf{core}_X(K)$. It follows that $N < K$. Thus, $K/N$ acts regularly on $\Gamma_N$. This contradicts that the Petersen graph has no vertex regular subgroup. Hence, $\Gamma$ is not a Cayley graph.
\end{proof}

 \begin{lemma}
  The graph $\Gamma$ has girth 10.
 \end{lemma}
 \begin{proof}
 We denote the images of elements, subgroups, or subsets in $X$ under the natural homomorphism to $X/N$ by placing a bar over them.
 It is convenient to use the same symbol to denote the graph homomorphism from $\Gamma$ to $\Gamma_N$, which maps each vertex $Hx \in V\Gamma$ to $\overline{Hx}=\overline{H}\overline{x}=\braket{(123),(12)(45)} \overline{x}$.

    Set $g_1=g, g_2=g^{h_1}$ and $g_3=g^{h_1^2}$. It is easily seen that $HgH$ is a disjoint union of $Hg_i$ for $1\leq i\leq 3$. Thus, for any vertex $Hx\in V\Gamma$, the neighbourhood of $Hx$ is $\{Hg_ix| 1\leq i\leq 3\}$. It follows that $(v_0, v_1,\dots, v_s)$ is an $s$-arc of $\Gamma$, if and only if there is an element $x\in G$ and a sequence $(y_1, y_2,\dots, y_s)$ where $y_i\in \{g_1, g_2, g_3\}$ such that
    \begin{align*}
    v_0&=Hx,& v_1&=Hy_1x, &&\dots &v_i&=Hy_iy_{i-1}\dots y_1x,&& \dots  &v_s&=Hy_sy_{s-1}\dots y_1x.
    \end{align*}
    and $y_i\neq y_{i+1}$ for $1\leq i \leq s-1$.
    Note that an $s$-arc  $(v_0, v_1,\dots, v_s)$  is a cycle of length $s$ if and only if $v_s=v_0$ and $v_0, v_1,\dots, v_{s-1}$ are pairwise distinct.

    We first construct a $10$-cycle in $\Gamma$, which implies that the girth of $\Gamma$ is less than or equal to $10$.
    By calculation, we have $\braket{g_1,g_2}=\braket{g_1g_2}{:}\braket{g_1}\cong D_{10}$ and $\braket{g_1,g_2}\cap \braket{h_1,h_2}=1$.
    Now, set $y_{2i-1}=g_2, y_{2i}=g_1$ for $1\leq i\leq 5$, and let $v_0=H$, and $v_i=Hy_iy_{i-1}\dots y_1(1\leq i \leq 10)$ be vertices in $\Gamma$. Then
    \[v_{10}=H(g_2g_1)^5=H((g_1g_2)^{-1})^5=H((g_1g_2)^{5})^{-1}=H=v_0.\]
    and $(v_0,v_1, \dots, v_{10})$ is a $10$-arc in $\Gamma$.
    If there are two integers $0\le i<j\le 9$ such that $v_i=v_j$, then we have $Hy_jy_{j-1}\dots y_{i+1}=H$. That is $x=y_jy_{j-1}\dots y_{i+1}$ is an element in  $\braket{g_1,g_2} \cap \braket{h_1,h_2}=1$. Hence $x=1$. As $0\le i<j\le 9$ and $y_{i+1}, \dots , y_j\in \{g_1, g_2\}$, this is impossible. Thence, $v_0, \dots, v_9$ are pairwise distinct, and $(v_0,v_1, \dots, v_{10})$ gives rise to a $10$-cycle in $\Gamma$.

    Now suppose that $\Gamma$ is of girth $s$ for some integer $3\leq s\leq 9$ and let  $(v_0, v_1, \dots, v_s)$ be an $s$-cycle in $\Gamma$. By the $2$-arc-transitivity of $\Gamma$, without loss of generality, we may assume $v_0=v_s=H$, $v_1=Hg_1$, $v_2=Hg_2g_1$. Moreover,
    there is a subscript sequence $(i_1, i_2, \dots, i_s)$,
    \begin{equation}\label{eqns1}
    i_j\in \{1,2,3\}(1\le j\le s),~ i_1=1,~ i_2=2,~\mbox{and}~ i_j\ne i_{j-1}(3\le j\le s),
    \end{equation}
    such that $v_j=Hg_{i_j}g_{i_{j-1}}\dots g_{i_1}(1\le j\le s)$.
    As $H=v_0=v_s=Hy_sy_{s-1}\dots y_1$,  $g_{i_1}g_{i_2}g_{i_3}\dots g_{i_s}=(g_{i_s}g_{i_{s-1}}\dots g_{i_1})^{-1}\in H$.
    This gives
    \begin{equation}\label{eqns2}
     \overline{g_{i_1}}\cdot \overline{g_{i_2}}\cdot \overline{g_{i_3}} \dots \overline{g_{i_s}} =\overline{g_{i_1}g_{i_2}g_{i_3}\dots g_{i_s}}
    \in \overline H=\braket{(123),(12)(45)}.
    \end{equation}
    Using MAGMA \cite{MAGMA}, it is easy to find all possible sequences $(i_1, i_2, \dots, i_s)$ satisfying conditions~(\ref{eqns1}) and (\ref{eqns2}) for $3\leq s\leq 9$.
    See the first column in Table~\ref{tab:list}. Further calculations allow us to uniquely rewrite the element $g_{i_1}g_{i_2}g_{i_3}\dots g_{i_s}$ in the form $nh$, where $n\in N$ and $h\in H$.  See the second and the third columns in Table~\ref{tab:list}. As $g_{i_1}g_{i_2}g_{i_3}\dots g_{i_s}\in H$, we have $n=1$.
   Note that $P=\braket{\mu, \lambda}$ is perfect group, and therefore the elements
     \[\mu, \lambda\mu\lambda^{-1}, [\lambda, \mu], [\mu, \lambda], [\mu, \lambda^{-1}]\]
     in $P$ are not the identity element.
Also, since
     \[P=\braket{\lambda,\lambda^\mu}=\braket{\lambda,(\lambda^{-1})^\mu}=\braket{\lambda^{-1},(\lambda^{-1})^\mu}\]
   we have the commutator $[\lambda,\lambda^\mu]=\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda\mu$, which is not the identity in $P$. Similarly,
    \[ [\lambda,(\lambda^{-1})^\mu]=\lambda^{-1}\mu\lambda\mu\lambda\mu\lambda^{-1}\mu~\mbox{and}~[(\lambda^{-1})^\mu,\lambda^{-1}]=\mu\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}.\]
    are also not the identity.
    It follows that the framed entries in Table~\ref{tab:list} are not identity. Thus, for each row in Table~\ref{tab:list}, we all have $n\neq 1$, a contradiction.
     Hence, the shortest cycle in $\Gamma$ is of length $10$, and $\Gamma$ is of girth $10$.
 \end{proof}
		\begin{table}[ht]
	\scriptsize
		\begin{tabular}{c| c c }
		
		\hline
			$(i_1, i_2, \dots, i_s) $& $n$ & $h$\\ \hline
    $( 1, 2, 1, 2, 1 )$&  $(\mu^\lambda , \qd{\mu},\mu^{\lambda\mu},\mu^\lambda,\mu)$   & $ h_2h_1$    \\
    $( 1, 2, 3, 1, 2 )$&$  (\qd{[\lambda,\mu]}, [\mu,\lambda^{-1}], [\mu,\lambda^{-1}], [\mu,\lambda^{-1}], [\mu, \lambda] )  $&$  h_1^2   $\\ \hline
    $( 1, 2, 1, 3, 1, 2 )$&$  (\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda^{-1}\mu,\qd{\mu},\lambda\mu\lambda^{-1},\mu\lambda\mu\lambda\mu\lambda)  $&$    h_2 $\\
    $( 1, 2, 3, 2, 1, 2 )$&$   (\qd{\mu},\mu\lambda\mu\lambda\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu,\mu,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}) $&$   h_1h_2  $\\ \hline
    $( 1, 2, 1, 3, 2, 1, 2 )$&$  (\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\mu\lambda\mu\lambda\mu\lambda\mu,\qd{[\mu, \lambda]},\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda^{-1})   $&$   h_1  $\\
    $( 1, 2, 3, 2, 3, 2, 1 )$&$  (\lambda\mu\lambda^{-1},\mu\lambda\mu\lambda^{-1}\mu,\qd{\mu},\mu\lambda^{-1}\mu\lambda\mu,\mu\lambda^{-1}\mu\lambda\mu)  $&$   h_2  $\\ \hline
    $( 1, 2, 1, 2, 3, 1, 3, 1 )$&$   (\lambda^{-1}\mu\lambda\mu\lambda^{-1}\mu\lambda\mu ,\qd{[\mu, \lambda]},\mu\lambda^{-1}\mu\lambda,\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\mu\lambda\mu\lambda^{-1})   $&$   h_1  $\\
    $( 1, 2, 1, 2, 3, 2, 1, 3 )$&$  (\lambda^{-1}\mu\lambda\mu\lambda\mu\lambda^{-1},\mu\lambda^{-1}\mu\lambda\mu,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\qd{\mu},\mu\lambda\mu\lambda^{-1}\mu)  $&$  h_1h_2   $\\
    $( 1, 2, 1, 3, 2, 3, 2, 1 )$&$   (\lambda^{-1}\mu\lambda\mu,\qd{[\mu,\lambda^{-1}]},\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu,\lambda\mu\lambda\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu\lambda\mu\lambda^{-1}) $&$  h_1^2   $\\
    $( 1, 2, 3, 1, 3, 1, 2, 3 )$&$ (\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\mu\lambda\mu\lambda^{-1}\mu\lambda\mu\lambda^{-1},\mu\lambda\mu\lambda\mu\lambda\mu,\mu\lambda^{-1}\mu\lambda,\qd{[\mu,\lambda^{-1}]})   $&$  h_1    $\\
    $( 1, 2, 3, 1, 3, 2, 3, 2 )$&$  (\lambda^{-1}\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu,\mu\lambda\mu\lambda\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu,\qd{\mu})    $&$  h_1^2h_2   $\\
    $( 1, 2, 3, 2, 3, 1, 2, 1 )$&$  (\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda\mu\lambda\mu,\mu\lambda^{-1}\mu\lambda,\qd{[\mu,\lambda]},\mu\lambda^{-1}\mu\lambda\mu\lambda^{-1}\mu\lambda)  $&$ h_1     $\\ \hline
    $( 1, 2, 1, 2, 3, 1, 2, 1, 3 )$&$( \qd{[\lambda,\mu]} ,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu,\lambda^{-1}\mu\lambda\mu,\mu\lambda\mu\lambda\mu\lambda\mu)    $&$   h_1^2   $\\
    $( 1, 2, 1, 2, 3, 2, 3, 1, 3 )$&$ (\qd{\lambda^{-1}\mu\lambda\mu\lambda\mu\lambda^{-1}\mu},1,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda,1,\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda)     $&$  1   $\\
    $( 1, 2, 1, 3, 1, 3, 1, 2, 3 )$&$  (\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda^{-1}\mu,\qd{\mu},\mu\lambda\mu\lambda^{-1}\mu)  	  $&$     h_1h_2 $\\
    $( 1, 2, 1, 3, 1, 3, 2, 3, 2 )$&$  (\qd{\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda\mu},\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda,1,1,\mu\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}) 	  $&$  1   $\\
    $( 1, 2, 1, 3, 2, 3, 1, 2, 1 )$&$  (\qd{\mu},\mu\lambda\mu\lambda\mu\lambda,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1},\mu\lambda\mu\lambda^{-1}\mu)    $&$  h_1h_2   $\\
    $( 1, 2, 3, 1, 3, 1, 3, 2, 3 )$&$   (\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda,\mu\lambda\mu\lambda^{-1}\mu\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda^{-1}\mu,\qd{\mu},\mu\lambda\mu\lambda^{-1}\mu) $&$   h_1h_2   $\\
    $( 1, 2, 3, 1, 3, 2, 1, 2, 3 )$&$   (\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda\mu\lambda,\qd{\lambda\mu\lambda^{-1}})   $&$  h_2   $\\
    $( 1, 2, 3, 2, 1, 3, 1, 2, 3 )$&$  (\mu\lambda\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}\mu,\mu\lambda\mu\lambda\mu\lambda,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1},\qd{\mu})    $&$    h_1^2h_2  $\\
    $( 1, 2, 3, 2, 1, 3, 2, 3, 2 )$&$   (\qd{[\mu,\lambda]},\mu\lambda\mu\lambda\mu\lambda\mu,\mu\lambda\mu\lambda\mu\lambda\mu,\mu\lambda^{-1}\mu\lambda,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda^{-1}\mu)   $&$    h_1   $\\
    $( 1, 2, 3, 2, 3, 1, 3, 1, 2 )$&$   (1,\qd{\mu\lambda\mu\lambda\mu\lambda^{-1}\mu\lambda^{-1}},1,\mu\lambda^{-1}\mu\lambda^{-1}\mu\lambda\mu\lambda,\mu\lambda^{-1}\mu\lambda\mu\lambda\mu\lambda^{-1})     $&$  1   $\\ \hline
		\end{tabular}

	\normalsize
		\vspace*{0.3cm}
		\caption{The subscript sequences $(i_1, i_2, \dots, i_s) $ and corresponding group elements $n\in N$ and $h\in H$ with $g_{i_1}g_{i_2}g_{i_3}\dots g_{i_s}=nh$.}
		\label{tab:list}
	\end{table}

\vskip0.2in
\thanks{{\bf Acknowledgements.}
This work was supported by the National Natural Science Foundation of China (No.1931005,12101518,11901512), the Fundamental Research Funds for the Central Universities (No. 20720210036, 20720240136), and the Natural Science Foundation of Yunnan Province (No. 202401AT070275). }

\bibliographystyle{abbrv}
\bibliography{reference}	
		
\end{document}